[-3+(-6)]SQAURED - [-8-2(-3)]SQAURED
I MULTIPLIED NEG 3 TO NEG 6 GOT 18. THEN SQAURED, GOT 81.
SECOND PART IS WHERE I GET STUCK-
NEG 8 NEG 2 = POS 10.THEN I TRIED MULTIPLYING TO 3SQUARED, IT WAS WRONG,( ACCORDING TO THE BACK OF THE BOOK) SO THEN I TRIED, 10 MINUS 9 GOT 1-WRONG.
THEN I MULTIPLIED 13 TO 13, GOT 169 MINUS 81-GOT 88. WRONG AGAIN.
I HAVE JUST ABOUT PUT A WHOLE THROUGH MY PAPER. PLEASE HELP. I HAVE A MAJOR TEST TOMORROW. ALSO, I WOULD LIKE TO KNOW WHY b-c, IF b=-7 AND c=12 ISN'T 5 BUT NEG 12?!?! THANK YOU

Answer by pwac(253)

in first bracket you are asked to add two-ve numbers
-3 + -6+=-9 which squared = 81
second bracket -8-2=-10 x -3=30 which squared = 900
so 81-900= -819 hope this is answer in book
In second question a negative number minus a positive number gives
a larger negative number.
Anyhow the figures in your problem do not work out
If b=-7 and c=12 then b-c would equal -19
Are sure c is not =5 ?
Pete